statistics. Healthy women aged 18 to 40 participated in a study of eating habits. Subjects were given bags of potato chips and bottled water and invited to snack freely. Interviews showed that some women were trying to restrain their diet out of concern about their weight. How much effect did these good intentions have on their eating habits? Here are the data on grams of potato chips consumed (note that the study report gave the standard error of the mean rather than the standard deviation): Group n x⎯⎯⎯ SEM Unrestrained 9 62.98 7.02 Restrained 11 29 10 The standard error of the difference of sample means (±0.0001) is: . The critical value (±0.001) from the t distribution for confidence interval 90% using the conservative degrees of freedom is: t* = . Give a 90% confidence interval (±0.01) that describes the effect of restraint: to . Based on this interval, is there a significant difference between the two groups?
Answer #1

standard error of difference

= sqrt( 7.02^2 + 10^2)

= 12.218

df = 18

critical value for 90% confidence level = 1.740

diff in means = (62.98 - 29) = 33.98

hence

confidence interval

(33.98 +- 1.740 * 12.218)

= ( 12.7207 , 55.2393 )

yes, because 0 is not present in confidence interval